### v3

# Charged pions asymmetry due to interference from the decay of light neutral axial mesons

Abstract
Effective three meson couplings, based in flavor U(3) quark-antiquark interactions, are considered to describe two light neutral axial meson Strong decays, $A\sim f_{1}(1285)$ and $f_{1S}(1420)$, and the charged rho meson decay in the following channels: $A\to\rho^{\mp}(770)+\pi^{\pm}$, $A\to a_{0}^{\mp}(980)+\pi^{\pm}$ and $\rho^{\pm}\to\pi^{\pm}\pi^{0}$. By considering neutral meson mixings, such as $f_{1}(1285)-a_{1}^{0}(1260)$, $f_{1S}(1420)-a_{1}^{0}(1260)$ and $\eta-\pi^{0}$ for the channels investigated, the leading three-meson interactions may lead to decay amplitudes that undergo interference and a resulting ratio of decay rate into charged pions, $\Gamma_{\pi^{+}}/\Gamma_{\pi^{-}}$, slightly smaller than one. For that it will be assumed that the produced charged vector mesons or charged scalar mesons undergo absorption, decay into channels without charged pions, or any other inelastic process that may suppress their decays into charged pions.
1 Introduction
Effective three-meson interactions may describe particular channels of meson (strong) decay into two mesons and several types have been derived and investigated with different approaches [1, 2, 3, 4, 5, 6, 7]. These couplings may correspond to a mechanism for meson production in (relativistic) heavy ion collisions, usually based in quark polarization processes. Meson production has been extensively investigated in the last decades, being their description usually difficult since they may have associated to final state strong interactions [8, 9, 10, 11, 12, 13]. Several facilities and experiments investigate and plan to investigate related aspects in a wide regime of energies, that may involve meson production, polarization and their decays, such as LHCb, BESIII, EIC, COMPASS, FAIR/GSI, J-PARC, JLAB, NICA, HERMES, SLAC. Asymmetry in charged pion production is an important observable in intermediary and high energies processes in heavy ions collisions, for example in [8, 9]. Several different physical effects from finite density environment may contribute to the ratio $\pi^{-}/\pi^{+}$ close to three [12, 13, 14] being that it decreases with the beam energy (with $A$ GeV), and it increases with the $N/Z$ asymmetry [11]. This ratio is related to the dynamics and behavior of hadrons under different conditions and it helps to determine freeze-out curves [10]. There is also an intense and wide investigation in sources of CP breaking at high energies by analyzing weak interactions in heavy meson decays, few examples in [15, 16, 17, 18, 19]. In the present work, the possibility of CP breaking in decays of light axial meson is envisaged by means of interference effects.
Three meson interactions based in QCD- one loop quark polarization have been calculated by considering, for example, the Global Color Model (GCM) or the Nambu-Jona-Lasinio (NJL)-model [20, 21, 22, 6, 7]. Dynamical Chiral Symmetry Breaking (DChSB) endows constituent quarks with large quark masses and this makes possible large quark mass expansions of the quark determinant. Flavor-Dirac quark currents provide suitable states to introduce meson multiplets, already coping with the constituent quark model, what can be implemented by means of the auxiliary field method. However, there are longstanding controversial results concerning light scalar and axial mesons structures, for which non-quark-antiquark states may contribute considerably [23, 25, 24]. Nevertheless, along this work, it will be assumed that an axial meson nonet and a light scalar meson nonet may be defined, although the physical states may be a result of superposition with different non-quark-antiquark components. In particular, the following scalar and axial mesons will be considered as part of flavor U(3) nonets: the isotriplets scalar $a_{0}(980)$ and axial $a_{1}(1260)$, and isoscalar axial mesons $f_{1}(1285)$ and $f_{1S}(1420)$ [25]. The production of axial $f_{1}(1285)$ and $f_{1S}(1420)$ has been investigated in proton-proton collisions [26, 27, 28] and it may involve different types of processes. In the present work, the decay of these two axial mesons into charged pions and another meson - either a charged $\rho(770)$ or charged scalar $a_{0}(980)$, will be investigated with three-meson vertices.
Therefore, in this work, some specific three light meson leading interactions are re-derived, along the lines of [7] and used to investigate neutral axial meson decays into charged pions. These flavor U(3) interactions describe several channels, among which only processes involving mesons composed exclusively by up and down quarks in the final states will be considered. The channels with strange mesons will necessarily involve states for which there are still further problems to define quark-antiquark mesons, possibly in possible scalar and axial meson nonets, and they will be investigated in a separate work. Furthermore, the point-like limit of the three-meson interactions, and the corresponding zero momentum limit of the form factors, will be considered. This will provide an upper bound for the resulting effects. The following decays of neutral axial mesons will be considered ($A=f_{1}(1285)$ and $f_{1S}(1420)$): $A\to\pi^{\mp}+a_{0}^{\pm}$ (processes containing one axial, one scalar and one pseudoscalar mesons will be labeled by ASP) and $A\to\pi^{\mp}+\rho^{\pm}$ (processes containing one axial, one vector and one pseudoscalar mesons will be labeled by VPA). To complement the description, the decay $\rho^{\pm}\to\pi^{\pm}+\pi^{0}$ with large relative branching ratio - will also be presented. The work is organized as follows. In the next section, the flavor U(3) quark determinant with background (constituent) quark currents is considered in the limit of local quark currents that give rise to the corresponding light meson multiplets by means of the auxiliary field method (AFM). All the local couplings are resolved by means of a zeroth order derivative expansion in the limit of zero momentum exchange, leading to coupling constants along the lines presented in [7]. In section (2.1) three-meson interactions for two channels of neutral axial meson decays, $f_{1}(1285)$ and $f_{1S}(1420)$, into charged pions and rhos $\rho^{\pm}(770)$ or into charged pions and scalars $a_{0}^{\pm}(980)$ will be presented, named respectively as VPA and ASP. For the channel VPA a next leading interaction will be also considered as shown in section (2.1.1). In section (2.2) the leading interactions describing the decay $\rho^{\pm}\to\pi^{\pm}\pi^{0}$, VPP, are presented. In sections (2.3) and (2.4) the renormalization conditions fixing free parameters of the AFM and the definition of meson states, considered along the work, are presented. In section (3) the different ratios $\Gamma_{\pi^{+}}/\Gamma_{\pi^{-}}$ for each of the decays are calculated by including meson mixings. Numerical results are exhibited in (4). In the last section there is a summary.
2 Quark determinant with background quark currents: and meson local interactions
The quark determinant in the presence of background constituent quark currents can be written as [6, 30, 7]:
$S_{eff}$ $=$ $-i\;Tr\;\ln\;\left\{i\left[{S_{0}}^{-1}+\sum_{\phi}a_{\phi}J_{\phi}\right]\right\},$ (1)
where $Tr$ stands for traces of all discrete internal indices and integration of space-time coordinates, $S_{0}$ is the quark propagator with a constituent quark mass due to the DChSB ($M$). The leading background dressed quark currents selected to contribute are the following:
$\sum_{\phi}a_{\phi}J_{\phi}=K_{0}\left[2R(x-y)(J_{S}^{i}(x,y)+i\gamma_{5}\lambda_{i}J_{PS}^{i}(x,y))-{R}^{\mu\nu}(x-y)\gamma_{\mu}\lambda_{i}(\gamma_{5}J_{\mu}^{V,i}(x,y)+\gamma_{5}J_{\mu}^{A,i}(x,y))\right],$ (2)
where $\lambda_{i}$ are the flavor Gell Mann matrices, flavor indices of the adjoint representation are $i,j=0...(N_{f}-1)^{2}$, $K_{0}=2\alpha_{g}^{2}/9$ being $2/9$ from a Fierz transformation and $\alpha_{g}$ the running quark-gluon coupling constant. The quark currents were defined respectively as $J_{S}^{i}=J_{S}^{i}(x,y)=(\bar{\psi}\lambda^{i}\psi)$, $J_{PS}^{i}=J_{PS}^{i}(x,y)=(\bar{\psi}i\gamma_{5}\lambda^{i}\psi)$, $J_{V,i}^{\mu}=J_{V,i}^{\mu}(x,y)=(\bar{\psi}\gamma^{\mu}\lambda^{i}\psi)$ and $J_{A,i}^{\mu}=J_{A,i}^{\mu}(x,y)=(\bar{\psi}\gamma^{\mu}\gamma_{5}\lambda^{i}\psi)$. Note that, in Eq. (1), the background quark currents are dressed by components of the effective gluon propagator $R(x-y)$ and $R^{\mu\nu}(x-y)$ that are defined below, and they can be considered as constituent quark currents. The large quark mass expansion done below is also compatible with a large gluon effective mass expansion since all the coefficients $a_{\phi}$ in Eq. (1) contain components of the gluon propagator making the overall quantities $a_{\phi}J_{\phi}$ with relative small strength. Although the resulting model is non-local, the large quark mass (and to some extent the large effective gluon mass) regime can be considered suitable to take the local limit of such quark currents. In this work, the indices of the currents will be treated such as it does not matter whether they stand as superscripts or subscripts. The functions $R(x-y)$ and $R^{\mu\nu}(x-y)$ are expressed in terms of the longitudinal and transverse components of the gluon propagator and they can be written in momentum space as:
$R(k)$ $=$ $3R_{T}(k)+R_{L}(k),$
$R^{\mu\nu}(k)$ $=$ $g^{\mu\nu}(R_{T}(k)+R_{L}(k))+2\frac{k^{\mu}k^{\nu}}{k^{2}}(R_{T}(k)-R_{L}(k)).$ (3)
[…]
The auxiliary field method is extremely useful to implement quark-antiquark states from the quark currents. It produces quark model meson multiplets naturally based in flavor symmetry. However, instead of introducing the non-local auxiliary (meson) fields, the local limit of the above quark currents will be considered and local meson fields for each of the channels (scalar, pseudoscalar, vector and axial) will be introduced by means of functional delta functions [31] that can be written as
11 11 $=$ $N\int D[S_{i},P_{i}]\delta(S_{i}-G_{0,s}^{ii}J_{i})\delta(P_{i}-G_{0,p}^{ii}J_{P,i})\;\int D[V_{\mu},A_{\mu}]\delta(V^{\mu}_{i}-G_{0,v}^{ii}J_{V,i}^{\mu})\delta(A_{i}^{\mu}-G_{0,a}^{ii}J_{A,i}^{\mu}),$ (4)
where $N$ is a normalization, $S_{i},P_{i},V_{i}^{\mu},A_{i}^{\mu}$ are the meson fields of each of the flavor multiplet, $D[S_{i},P_{i}]$ and $D[V_{\mu},A_{\mu}]$ are the measures of integration, and $G_{0,s}^{ii},G_{0,p}^{ii},G_{0,v}^{ii}$ and $G_{0,a}^{ii}$ are constant parameters (dimension $M^{-2}$). These parameters contain the meson renormalization constants that are usually introduced with functional delta function such as the following: $\delta(\phi\sqrt{Z}_{\phi}-G_{0}^{ii}J_{i}^{\phi})$, where $\phi$ is a local meson fields and $G_{0}$ is a dimensionful constant [33, 32]. By adopting this latter form, the former expression (4) is recovered with redefinition of the normalization constant and of $G_{0}^{ii}$.
The third order large quark mass expansion of the above determinant produces three-meson interactions. The leading possible channels of interaction have been presented, for the local limit, in [7] by considering only the non-derivative or lowest order derivative terms. In the present work, four types of leading order (in derivatives) vertices containing pseudoscalar field mesons will be selected with an additional non-leading (NL) vertex for one of these leading ones. The coupling constants will be resolved within the zero order derivative expansion as presented in [34].
2.1 VPA and ASP vertices: neutral axial mesons decays
There are two types of leading vertices that yield an axial meson decay into a final state with only up and down quark mesons. The first has no spatial derivative acting on the fields and contains axial-vector-pseudoscalar (VPA) local meson fields. The second one has one spatial derivative acting in one of the fields and contains axial-scalar-pseudoscalar (ASP) mesons. In the local limit the following interactions are obtained:
${\cal L}_{6,sb,VPA}^{mes}$ $=$ $T^{ijk}\frac{3i\;G_{sb1}\left(V^{\mu}_{i}A_{\mu}^{j}P_{k}-A^{\mu}_{i}V_{\mu}^{j}P_{k}\right)}{G_{0,v}G_{0,ps}G_{0,a}},$ (5)
${\cal L}_{6d,ASP}^{mes}$ $=$ $3\;\frac{T^{ijk}}{G_{0,a}G_{0,s}G_{0,ps}}\left\{\left(\partial_{\mu}P_{i}\right)\left[-G_{d1}S_{j}A^{\mu}_{k}+G_{d2}A^{\mu}_{j}S_{k}\right]-G_{d2}(\partial_{\mu}S_{i})\left[A_{j}^{\mu}P_{k}+P_{j}A^{\mu}_{k}\right]\right.$ (6)
$+$ $\left.(\partial_{\mu}A^{\mu}_{i})\left[G_{d1}S_{j}P_{k}+G_{d2}P_{j}S_{k}\right]\right\},$
where $T^{ijk}=2(d_{ijk}+if_{ijk})$, in terms of the flavor group symmetric and anti-symmetric structure constants respectively. By defining the function $R^{\eta}(k)$ to be either the longitudinal $\eta=L$ or the transverse $\eta=T$ part of the gluon propagator, the following flavor dependent coupling constants were defined, for the limit of zero momentum exchange:
$G_{sb1}T_{ijk}\;g^{\mu\rho}$ $=$ $\;A_{2\eta}\frac{2N_{c}}{3}K_{0}^{3}Tr_{D,F}\int_{k}{S}_{0}(k)\gamma^{\mu}\lambda_{i}R^{\eta}(-k){S}_{0}(k)\gamma^{\rho}\lambda_{j}R^{\eta}(-k){S}_{0}(k)\lambda_{k}R^{\eta}(k),$
$G_{d1}T_{ijk}g^{\mu\rho}$ $=$ $A_{1\eta}\frac{4N_{c}}{3}K_{0}^{3}Tr_{D,F}\int_{k}\tilde{S}_{0}(k)\gamma^{\rho}\gamma^{\mu}\gamma_{5}\lambda_{i}R^{\eta}(-k){S}_{0}(k)\lambda_{j}R^{\eta}(-k){S}_{0}(k)\lambda_{k}\gamma_{5}{R}^{\eta}(k),$
$G_{d2}T_{ijk}g^{\mu\rho}$ $=$ $A_{1\eta}\frac{4N_{c}}{3}K_{0}^{3}Tr_{D,F}\int_{k}\tilde{S}_{0}(k)\gamma^{\rho}\gamma^{\mu}\gamma_{5}\lambda_{i}R^{\eta}(-k){S}_{0}(k)\gamma_{5}\lambda_{j}R^{\eta}(-k){S}_{0}(k)\lambda_{k}{R}^{\eta}(k),,$ (7)
[…]
$A_{1T}=\frac{3}{2},\;\;\;\;A_{1L}=\frac{1}{2},\;\;\;A_{2T}=3,\;\;\;A_{2L}=1.\;\;\;$ (8)
It is interesting to note that, in general, the coupling constants for a particular two meson decay channel are the same for both types of interactions with symmetric and antisymmetric flavor structure constants $d_{ijk}$ and $f_{ijk}$ This means that the specific momentum dependence of such form factors, defined by Eqs. (2.1) and in other equations below, should nearly factorize in the ratios of decay rates explored below. Therefore, these momentum dependencies should not introduce considerable changes in the resulting ratios of meson decay, and the coupling constants will be considered for numerical calculations instead of the form factors. Furthermore, this issue must be related to the problem of fixing the vector and axial meson polarizations for all the corresponding momentum dependent vertices. The inclusion of the full momentum dependent form factors and their behavior, conjugated to the choice of the vector/axial meson polarization, as discussed after Eq. (3.1) and below, will be developed in another work. After having defined the coupling constants at the zero momentum exchange, i.e. neglecting the momentum dependence, all the energy dependencies of the couplings in Eq. (6) will also be neglected by fixing the longitudinal vector and axial meson polarizations, such as $\mu=0$, and fixing the meson energy as its rest energy, i.e. $E_{\phi}=M_{\phi}$.
All the coupling constants are ultraviolet finite. Although the resulting effective model is non-renormazible, effects of these coupling constants of meson dynamics make sense by means of a suitable renormalization to fix the parameters $G_{0}^{ii}$, what is discussed below. The coupling constant $G_{sb1}$ is proportional to the quark mass that has been taken as constant. Therefore, it corresponds to (chiral) symmetry breaking term, and it disappears in the limit of massless quarks. $G_{d1},G_{d2}$ are non-zero, even in the case of massless quarks. Whereas in the ASP channel there are terms of both types with $f_{ijk}$ and $d_{ijk}$, for the channel VPA only the antisymmetric structure constant contributes. Therefore, a next-leading (NLO) VPA term will also be taken into account.
2.1.1 NLO VPA
One next leading higher order derivative term will be considered now that will be referred to as ddVPA. There are several contributions that, in the local limit, can be written as:
${\cal L}_{ddVPA}$ $=$ $3\;i\;\frac{T_{ijk}G_{sva,sb}}{G_{0,p}G_{0,v}G_{0,a}}\left\{\Gamma^{\nu\alpha\beta\mu}A_{\nu}^{i}(\partial_{\alpha}P_{j})(\partial_{\beta}V_{\mu}^{k})+\Gamma^{\nu\beta\mu\alpha}A_{\nu}^{i}(\partial_{\beta}V_{\mu}^{j})(\partial_{\alpha}P_{k})-\Gamma^{\alpha\nu\beta\mu}(\partial_{\alpha}A_{\nu}^{i})P_{j}(\partial_{\beta}V_{\mu}^{k})\right.$ (9)
-- $\left.\Gamma^{\alpha\nu\beta\mu}(\partial_{\alpha}A_{\nu}^{i})(\partial_{\beta}V_{\mu}^{j})P_{k}+\Gamma^{\beta\nu\alpha\mu}(\partial_{\beta}A_{\nu}^{i})(\partial_{\alpha}P_{j})V_{\mu}^{k}-\Gamma^{\alpha\beta\nu\mu}(\partial_{\alpha}P_{i})(\partial_{\beta}A_{\nu}^{j})V_{\mu}^{k}\right\},$
where the following tensor was defined:
$\Gamma^{\mu\alpha\beta\rho}$ $=$ $\left(g^{\mu\alpha}g^{\rho\beta}-g^{\alpha\rho}g^{\mu\beta}+g^{\beta\alpha}g^{\mu\rho}\right),$ (10)
and the coupling constant, for the limit of zero momentum exchange, was defined as:
$G_{sva,sb}T_{ijk}\Gamma^{\mu\alpha\beta\rho}$ $=$ $\frac{2N_{c}}{3}K_{0}^{3}\;Tr_{D,F}\int_{k}{S}_{0}(k)\gamma_{\nu}\gamma^{5}\lambda_{i}R^{\mu\nu}(k)\tilde{S}_{0}(k)\gamma^{\alpha}i\gamma_{5}\lambda_{j}R(-k)\tilde{S}_{0}(k)\gamma^{\alpha}i\gamma_{\sigma}\lambda_{k}R^{\rho\sigma}(-k),$ (11)
To deal with such large number of terms, they were rewritten by performing an integration by parts to eliminate the derivative $\partial_{\alpha}V_{\mu}$. Besides that, the same remarks written after Eq. (8) applies: vector and axial meson polarizations were chosen to be the longitudinal, $V_{\mu}\to V_{0}$ and $A_{\mu}\to A_{0}$. As a consequence, most of the terms depend directly on the pion and axial meson energies ($E_{\pi}$ and $E_{A}$). However, since the momentum exchange of the couplings was neglected in their integrals, the energy of the mesons will be fixed as their rest masses. With these choices, the vertices obtained from the above equation can be written as:
${\cal V}_{ddVPA}$ $=$ $-6\;i\;\frac{G_{sva,sb}}{G_{0,p}G_{0,v}G_{0,a}}\left\{2d_{ijk}A_{0}^{i}V_{0}^{j}P^{k}(E_{\pi}E_{A}+E_{\pi}^{2}-E_{A}^{2})+2if_{ijk}A_{0}^{i}V_{0}^{j}P^{k}(E_{\pi}E_{A})\right\}.$ (12)
2.2 VPP: Charged rho meson decay into two pions
In this section we develop vertices with one vector meson and two pseudoscalar mesons, labeled by VPP. There are several ways to treat the quark determinant. Strictly, the expansion shown above does not describe several channels, such as the charged vector meson decays into two pions, $\rho^{\pm}\to\pi^{\pm}\pi^{0}$. To do that, let us define the following quantity in the determinant
$X=S_{0}(k)\;K_{0}\;R^{T,L}(k)\;J_{\phi}$ (13)
Eq. (1) will be worked out, except for an irrelevant constant, as
$-i\;Tr\;\ln\;\left\{i[1+X]\right\}=-\frac{i}{2}Tr\;\ln\;\left\{[1+X][1+X^{*}]\right\}.$ (14)
The local auxiliary fields can then be introduced in the same way by reducing the currents to their local limit after resolving the coupling constants. In the second order expansion, mass terms of charge (and mass) eigenstates as combinations of the form:
${\cal L}$ $=$ $-\frac{1}{2}\sum_{\phi}M_{\phi_{i}}^{2}\phi_{i}^{2}\cdot{\phi_{i}^{*}},\;\;\;\;\;\phi_{i}=S_{i},P_{i},V_{\mu,i},A_{\mu,i},$ (15)
where $M_{\phi_{i}}^{2}$ can be calculated straightforwardly, but will not be developed in the present work. To obtain interactions that describe explicitly the desired decay, three third order terms with a single pseudoscalar complex conjugate field will be selected with the single derivative acting in one of them such as to compose a positive parity Lorentz scalar interaction term. The terms with such structure, in the local limit, can be written as:
${\cal L}_{VPP}^{mes}$ $=$ $3T^{ijk}\left[iG_{d2}\left(\left\{(\partial^{\mu}V_{\mu}^{i})P_{j}^{*}P_{k}\right\}-\left\{V_{\mu}^{i}(\partial^{\mu}P_{j}^{*})P_{k}\right)\right\}+iG_{d1}\left\{V_{\mu}^{i}P_{j}^{*}(\partial^{\mu}P_{k})\right\}\right].$ (16)
In this expression, both types of interactions, with the symmetric and with the antisymmetric structure constants, will contribute to the decay $\rho^{\pm}\to\pi^{\pm}\pi^{0}$.
2.3 Axial meson mixings and renormalization conditions
The parameters $G_{0\phi}$ introduced with the local auxiliary meson fields, Eq. (4), can be fixed by a suitable renormalization of the meson fields, along the lines presented in [33, 6, 7]. For that, consider the following second order meson kinetic terms obtained from the determinant expansion:
${\cal L}_{kin}$ $=$ $\frac{1}{2}\frac{I_{0,S}^{ii}}{(G_{0,S}^{ii})^{2}}\partial_{\mu}S_{i}\partial^{\mu}S_{i}+\frac{1}{2}\frac{I_{0,S}^{ii}}{(G_{0,P}^{ii})^{2}}\partial_{\mu}P_{i}\partial^{\mu}P_{i}-\frac{1}{8}\frac{I_{0,V}^{ii}}{(G_{0,v}^{ii})^{2}}{\cal F}_{i}^{\mu\nu}{\cal F}_{\mu\nu}^{i}-\frac{1}{8}\frac{I_{0,A}^{ii}}{(G_{0,a}^{ii})^{2}}{\cal G}_{i}^{\mu\nu}{\cal G}_{\mu\nu}^{i},$ (17)
where the Abelian vector and axial meson tensors were defined as
${\cal F}_{\mu\nu}^{i}=\partial_{\mu}V_{\nu}^{i}-\partial_{\nu}V_{\mu}^{i},\;\;\;\;\;\;\;{\cal G}_{\mu\nu}=\partial_{\mu}A_{\nu}^{i}-\partial_{\nu}A_{\mu}^{i}.$
Some of the neutral meson fields develop mixing interactions totally analogous to the vector or pseudoscalar meson mixings shown in [33, 35, 36]. The following mixing terms will be considered:
${\cal L}_{mix}$ $=$ $\frac{I_{i\neq j}^{0,\phi}}{G_{0,\phi}^{ii}G_{0,\phi}^{jj}}\phi^{i}\phi_{j}\;\;\;\;\;i,j=0,3,8,\;\;\;\;\;\phi=S,P,V,A.$ (18)
Although energy dependent vector and axial meson mixings also arise, i.e. $\sim{\cal G}_{i}^{\mu\nu}{\cal G}_{\mu\nu}^{j}$, they are equivalent to the former by a field redefinition. In the present work, the energy independent limit of the interactions was adopted, since the coupling constants were defined in the zero momentum exchange limit.
The above equations fix the parameters $G_{0,\phi}$ by means of an ultraviolet cutoff adopted for the quantities $I_{0,\phi}^{ii}$. These terms give rise to the canonical field normalization terms, and this procedure settles the parameters $G_{0}^{ii}$. Consequently, renormalized three-meson coupling constants, by omitting the flavor indices ii that are all equal for the case of degenerate quark mass, can be defined, such as:
$G_{sb1}^{R}=\frac{G_{sb1}}{G_{0,v}G_{0,p}G_{0,a}},\;\;\;\;\;G_{d1/2}^{R}=\frac{G_{d1/2}}{G_{0,s}G_{0,p}G_{0,a}},\;\;\;\;\;G_{sva,sb}^{R}=\frac{G_{sva,sb}}{G_{0,v}G_{0,p}G_{0,a}}.$ (19)
The renormalized mixing parameters, with dimension $M^{2}$, can be written as:
$G^{R}_{i\neq j}$ $=$ $\frac{I_{0,\phi}^{ij}}{G_{0,\phi}^{ii}G_{0,\phi}^{jj}},\;\;\;\;\;i,j=0,3,8.$ (20)
Below, the mixing will be considered for the axial meson sector and, in a particular case, for the pseudoscalar $\pi^{0}-\eta$.
2.4 Meson states
[…]
$\frac{P_{1}\pm iP_{2}}{\sqrt{2}}\sim\pi^{\pm}(140),$ $P_{3}\sim\pi^{0}(135),$
$\frac{V_{1}\pm iV_{2}}{\sqrt{2}}\sim\rho^{\pm}(770),$ $V_{3}\sim\rho^{0}(770),\;\;\;\;\;\;\Gamma=0.147\;GeV$
$\frac{A_{1}\pm iA_{2}}{\sqrt{2}}\sim a_{1}^{\pm}(1260),$ $A_{3}\sim a_{1}^{0}(1260),\;\;\;\;\;\;\Gamma=0.420\;GeV,$
$\frac{S_{1}\pm iS_{2}}{\sqrt{2}}\sim a_{0}^{\pm}(980),$ $S_{0}\sim a_{0}(980),\;\;\;\;\;\;\Gamma=0.075\;GeV.$
$S_{0}\sim\eta(548)$ $M_{\eta}=548\;GeV,\;\;\;\;\;\Gamma_{\eta}=1.3\times 10^{-6}\;GeV.$ (21)
Besides that, the following states for the corresponding neutral axial mesons, with their masses and widths, will be considered below [25]:
$A_{0}$ $\sim$ $f_{1}(1285),\;\;\;M_{f_{1}}=1.252\;GeV,\;\;\;\;\Gamma=0.023\;GeV,$
$A_{8}$ $\sim$ $f_{1S}(1420),\;\;M_{f_{1S}}=1.426\;GeV,\;\;\;\Gamma=0.055\;GeV.$ (22)
where the widths were adopted as an average value of the range presented in PDG Table.
3 Asymmetries in the decay of light axial mesons into charged pions
In Figs. (1), (2) and (3) the specific processes that will be detailed and analyzed below are drawn: for the channels VPA, ASP and VPP. Figs. (1) and (3) show the decays of the two axial mesons $A_{0}$ and $A_{8}$ in different channels. The last one, Fig. (3), has the secondary decay of the rho into two pions, that is presented separately in Fig. (2). In the left-hand side (l.h.s.) columns of all these figures, there are diagrams obtained for the symmetric structure constant (d-channel) with square vertices. They correspond to two different interactions (ddVPA and ASP) of the neutral axial mesons $A_{0}\sim f_{1}(1285)$ and $A_{8}\sim f_{1S}(1420)$. In the right-hand side (l.h.s.) the amplitudes with the interactions defined with the antisymmetric structure constant (f-channel) with round vertices. Since the decays of axial mesons will be picked up among all the possible channels of the interactions VPA and ASP, the corresponding processes may be referred as VPA-A and ASP-A. The initial states considered below have heavier particles than the final state, and only energetically favorable decays will be considered. In some cases, due to the flavor channels involved, a meson mixing (diamond) is assumed to take place such that both d- and f-channels yield the same initial and final states. In Fig. (3) the rho meson decay vertices are indicated with a triangle that contains two components, d-channel and f-channel amplitude.
Figure 1: Three leg meson decays of two neutral axial mesons $A_{0}\sim f_{1}(1185)$ and $A_{8}\sim f_{1S}(1420)$, (2.4), with VPA and ddVPA interactions. There are d-channel (square) and f-channel (round) couplings involving charged pions. Intermediary states from axial meson mixings (diamonds) are considered in the antisymmetric channels.
Figure 2: Three leg meson decays of rho into two pions $\rho^{\pm}\to\pi^{\pm}\pi^{0}$. The rho decay is considered with a $\eta-\pi^{0}$ mixing, with diamond symbol, in the d-channel.
Figure 3: Three leg meson decays of two neutral axial mesons $A_{0}\sim f_{1}(1185)$ and $A_{8}\sim f_{1S}(1420)$ with VPA and ddVPA interactions. The d-channel (square) - lhs - and f-channel (round) - rhs - couplings involving charged pions. The charged rho is assumed to decay either into $\eta-\pi$ (d-channel) - is encoded in triangle symbol, and it includes both VPP d- and f- channels - shown in Fig. (2).
Most of the axial meson decays considered in this work have already been measured experimentally. The only decays for which no record was found in the PDG Table [25] is $a_{1}(1260)\to a_{0}^{0}(980)\pi$ and $f_{1S}(1420)\to\rho\pi$. The other processes, showed in figures (1), (2) and (3) are the following [25]:
[…]
Some of these decay may respect the OZI sum rule or G-parity conservation and others do not [24]. Besides that, neutral axial meson mixings might prevent a clear identification of their structure at the example of the $\eta-\eta^{\prime}$ mixing of the pseudoscalar mesons multiplet [43, 44]. Therefore, the channels presented above will be investigated below, at least, as a possible scenario. Besides that, there are decays of the charged rho $\rho$:
$\;\;\;\;\rho^{\pm}\to\pi^{\pm}\pi^{0},\;\;\;\;(\Gamma_{i}/\Gamma\sim 100\%),\;\;\;\;\;\;\;\;\;\;\;\;\rho^{\pm}\to\pi^{\pm}\eta,\;\;\;\;(\Gamma_{i}/\Gamma\sim 6\times 10^{-3}\%).$
3.1 VPA-A decay
The following decays will be analyzed in this section:
$A_{1}^{8}$$\to$ $\rho^{+}\pi^{-},\;\;\;\;\;A_{1}^{8}\to\rho^{-}\pi^{+},$
$A_{1}^{0}$$\to$ $\rho^{+}\pi^{-},\;\;\;\;\;A_{1}^{0}\to\rho^{-}\pi^{+}.$ (24)
Consider the antisymmetric part of the $VPA$ coupling, in Eq. (5), and of the ddVPA contribution, in Eq. (12), with $T_{ijk}\to 2if_{ijk}$. The vector and axial mesons will be considered in their longitudinal polarizations, $\mu=0$, such that they will be omitted. As discussed above, the energy dependence of the couplings will be reduced to the meson rest energies. By adding both contributions for the f-channel, it yields:
${\cal L}_{VPA}^{f}=6\;i\;f_{ijk}\left[2G_{sb1}^{R}+2E_{\pi}E_{A}\;G_{sva,sb}^{R}\right]V_{i}A^{j}P_{k},$ (25)
These terms can be written in terms of the corresponding mesons as charge/mass eigenstates with up and down quarks, i.e. for $i,j,k=1,2,3$. It yields:
${\cal L}_{A\rho^{\mp}\pi^{\pm}}^{f}=12\;i\;\left[G_{sb1}^{R}+E_{\pi}E_{A}\;G_{sva,sb}^{R}\right]\left[\pi_{3}(\rho^{-}a_{1}^{+}-\rho^{+}a_{1}^{-})+a_{1}^{0}(\rho^{+}\pi^{-}-\rho^{-}\pi^{+})+\rho_{3}(a_{1}^{-}\pi^{+}-a_{1}^{+}\pi^{-})\right],$ (26)
where $\pi,\rho$ and $a_{1}$ are components of the isotriplets: respectively pion, rho and axial $a_{1}$. The second interaction in this last equation corresponds to the r.h.s. of VPA-A in Fig. (1) before introducing the axial meson mixing that will be discussed below. The decay widths of the $f_{1}(1285)$ and of the $f_{1S}(1420)$ are not very large so that the mixing may be expected to occur. Note that there are different signs in the contribution of each of the positive and negative mesons for each coupling.
A related amplitude appears from the contributions of the vertex NLO-VPA (ddVPA) with the symmetric structure tensor, $d_{ijk}$. These terms can be written as:
${\cal L}_{VPA}^{d}$ $=$ $-12\;i\;d_{ijk}\;G_{sva,sb}^{R}(E_{A}E_{\pi}+E_{\pi}^{2}-E_{A}^{2})\;A^{i}V^{j}P^{k}.$ (27)
for which the following terms will be considered $d_{118}=d_{228}=\frac{1}{\sqrt{3}}$ and $d_{110}=d_{220}=\sqrt{\frac{2}{3}}$ respectively for axial mesons $A_{0}$ and $A_{8}$ decaying into charged pions. The processes of the l.h.s. of the first line of Fig. (1), correspond therefore to the following interactions:
${\cal L}_{A\rho^{\mp}\pi^{\pm}}^{d}$ $=$ $-12\;i\;G_{sva,sb}^{R}(E_{A}E_{\pi}+E_{\pi}^{2}-E_{A}^{2})\left[\frac{A_{8}}{\sqrt{3}}\left(\rho^{+}\pi^{-}+\rho^{-}\pi^{+}\right)+\sqrt{\frac{2}{3}}A_{0}\left(\rho^{+}\pi^{-}+\rho^{-}\pi^{+}\right)\right].$ (28)
3.1.1 Meson mixing
Consider the mixings between neutral meson flavor-eigenstates, obtained with meson auxiliary fields in the corresponding quark-antiquark interactions of the following form:
${\cal L}_{mix}=\sum_{ij}G_{ij}^{R}\phi_{i}\phi_{j},\;\;\;\;\;\;\;\;\;G_{i\neq j}\neq 0\;\;\mbox{for}\;\;i,j=0,3,8,$ (29)
The pseudoscalar and scalar mixing interactions $G_{i\neq j}$ have been calculated and used in [35, 36, 37]. These mixings give rise to to the neutral meson physical mass eigenstates $\varphi_{i}$ by means of a flavor rotation parameterized in terms of $\bar{G}_{\`{\i}\neq j}$ such that:
$\phi_{0}$ $\sim$ $\varphi_{0}+\bar{G}_{03}\varphi_{3}+\bar{G}_{08}\varphi_{8},$
$\phi_{3}$ $\sim$ $\varphi_{3}+\bar{G}_{30}\varphi_{0}+\bar{G}_{38}\varphi_{8},$
$\phi_{8}$ $\sim$ $\varphi_{8}+\bar{G}_{80}\varphi_{0}+\bar{G}_{83}\varphi_{3}.$ (30)
The neutral axial meson states, from $A^{(0,3,8)}$, can be approximated as:
$A_{8}$ $\sim$ $f_{1S}(1420)+\bar{G}_{38}a_{1}^{0}(1260)+\bar{G}_{80}f_{1}(1285),$
$A_{0}$ $\sim$ $f_{1}(1285)+\bar{G}_{30}a_{1}^{0}(1260)+\bar{G}_{80}f_{1S}(1420).$ (31)
These mixings can be applied in the Lagrangian terms of the form shown above:
${\cal L}_{mix}=\sum_{ij=0,3,8}G_{ij}^{R}A_{i}A_{j}.$ (32)
By requiring the new mixings among mass eigenstates $\varphi_{i}\varphi_{j\neq i}$ to disappear, and by picking up only each of the leading terms of each channel, the following meson mixing parameters are obtained:
$\bar{G}_{03}\simeq\frac{G_{03}^{R}}{2(G_{33}^{R}+G_{00}^{R})},\;\;\;\bar{G}_{08}\simeq\frac{G_{08}^{R}}{2(G_{88}^{R}+G_{00}^{R})},\;\;\;\bar{G}_{38}\simeq\frac{G_{38}^{R}}{2(G_{33}^{R}+G_{88}^{R})}.$ (33)
The resulting f-channel complete VPA-A amplitudes, from eq. (26), for the three leg vertices with neutral axial meson, with the mixing $A_{8}\to A_{3}$ and for $A_{0}\to A_{3}$, are the following:
${\cal V}_{A\rho^{+}\pi^{-}}^{f}$ ::: $12\;i\;\left[G_{sb1}^{R}+E_{\pi}E_{A}\;G_{sva,sb}^{R}\right]\;D_{A_{3}}(P^{2})\;\left(A_{1}^{8}\;\bar{G}_{38}+A^{0}_{1}\;\bar{G}_{03}\right),$
${\cal V}_{A\rho^{-}\pi^{+}}^{f}$ ::: $-12\;i\;\left[G_{sb1}^{R}+E_{\pi}E_{A}\;G_{sva,sb}^{R}\right]\;D_{A_{3}}(P^{2})\;\left(A_{1}^{8}\;\bar{G}_{38}+A^{0}_{1}\;\bar{G}_{03}\right).$ (34)
where $D_{A_{3}}(P^{2})$ is the intermediary state propagator, $A_{3}\sim a_{1}^{0}(1260)$, that will be written as a Breit Wigner propagator:
$D_{\phi}(P^{2})=\frac{i}{P^{2}-M_{\phi}^{2}+i\;M_{\phi}\;\Gamma_{\phi}},$ (35)
where $\phi\sim A_{3}\sim a_{1}(1260)$, in terms of its mass $M_{\phi}$ and decay width $\Gamma_{\phi}$.
The VPA amplitudes can be written in terms of the following quantities for the f- and d-channel:
$G_{A_{8}}^{f}(E_{A})=12\;\bar{G}_{38}\;\left(G_{sb1}^{R}+E_{\pi}E_{A}\;G_{sva,sb}^{R}\right),\;\;\;\;\;$ $G_{A_{0}}^{f}(E_{A})=12\;\bar{G}_{03}\;\left(G_{sb1}^{R}+E_{\pi}E_{A}\;G_{sva,sb}^{R}\right),$
$G_{A_{8}}^{d}(E_{A})=-\frac{12}{\sqrt{3}}\;G_{sva,sb}^{R}(E_{\pi}E_{A}+E_{\pi}^{2}-E_{A}^{2})\;,\;\;\;\;\;$ $G_{A_{0}}^{d}(E_{A})=-24\sqrt{\frac{2}{3}}\;G_{sva,sb}^{R}(E_{\pi}E_{A}+E_{\pi}^{2}-E_{A}^{2})\;.$
For the corresponding ratio of decay rates for the initial neutral axial vector with energy $P$ for each axial meson decay $A_{J}$ (J=0,8), and final states with positive or negative pions and rhos, one can write:
$\mbox{Ratio}\;(VPA-A_{J})\;:\;\frac{\Gamma_{\pi^{+}}^{J}}{\Gamma_{\pi^{-}}^{J}}$ $\sim$ $\frac{\left|+G_{A_{J}}^{d}+G_{A_{J}}^{f}D_{A_{3}}(P^{2})\right|^{2}}{\left|+G_{A_{J}}^{d}-G_{A_{J}}^{f}D_{A_{3}}(P^{2})\right|^{2}},\;\;\;\;\;\;\;\;\;\;J=0,8.$ (37)
3.2 ASP-A decay
The following ASP-A decays will be analyzed:
$A_{1}^{8}$$\to$ $a_{0}^{+}\pi^{-},\;\;\;\;\;A_{1}^{8}\to a_{0}^{-}\pi^{+},$
$A_{1}^{0}$$\to$ $a_{0}^{+}\pi^{-},\;\;\;\;\;A_{1}^{0}\to a_{0}^{-}\pi^{+}.$ (38)
By performing an integration by parts for the term with $\partial_{\mu}S_{k}$ in Eq. (6) all the derivative can be taken to act in the fields P or A. They can be written as:
${\cal L}_{6d,ASP}^{mes}$ $=$ $3\;T^{ijk}\left\{\left(\partial_{\mu}P_{i}\right)\left[(G_{d2}^{R}-G_{d1}^{R})S_{j}A^{\mu}_{k}+2G_{d2}A^{\mu}_{j}S_{k}\right]\right.$ (39)
$+$ $\left.(\partial_{\mu}A^{\mu}_{i})\left[(G_{d1}^{R}+G_{d2}^{R})S_{j}P_{k}+2G_{d2}^{R}P_{j}S_{k}\right]\right\}.$
The polarization of the axial meson will be longitudinal, $\mu=0$, and it will be omitted from here on. The term with the antisymmetric structure constant, with states containing only up and down quarks, involves $A_{3}\sim A_{1}^{(0)}$, the third component of the isotriplet $a_{1}(1260)$, similarly to the VPA channel. By selecting the channels involving charged pions, with $f_{123}=1$, and the neutral axial meson, it yields the following terms:
${\cal L}_{A_{1}\pi^{\pm}S^{\mp}}^{f}$ $=$ $6\;i\;G_{S^{f}}\;A_{1}^{3}\left[S^{-}(\pi^{+})-S^{+}(\pi^{-})\right],$ (40)
where
$G_{S^{f}}=G_{S^{f}}(E_{A})$ $=$ $\left[E_{\pi}(G_{d2}^{R}+G_{d1}^{R})-E_{A}(G_{d1}^{R}-G_{d2}^{R})\right].$ (41)
There are other channels involving pions, scalar and axial mesons that will not be analyzed. By considering the mixings $A_{1}^{8}-A_{1}^{3}$ and $A_{1}^{0}-A_{1}^{3}$ for the neutral axial meson, the corresponding amplitudes can be written as:
${\cal V}_{A\pi^{+}S^{-}}^{f}$ ::: $6\;i\;G_{S^{f}}\;D_{A_{3}}(P^{2})\left(\bar{G}_{38}A_{8}+\bar{G}_{03}A_{0}\right),$
${\cal V}_{A\pi^{-}S^{+}}^{f}$ ::: $-6\;i\;G_{S^{f}}\;D_{A_{3}}(P^{2})\left(\bar{G}_{38}A_{8}+\bar{G}_{03}A_{0}\right).$ (42)
From this expression the following quantities are defined:
$G_{S^{f}_{0}}$ $=$ $6\;G_{S^{f}}\bar{G}_{03},\;\;\;\;\;G_{S^{f}_{8}}=6\;G_{S^{f}}\bar{G}_{38}.$ (43)
For the symmetric interactions, since $d_{11n}=d_{22n}$ (n=0,8), one has the following amplitudes with charged pions and scalars and neutral axial meson:
${\cal V}_{A\pi^{\pm}S^{\mp}}^{d}$ $=$ $+\;i\;G_{S^{d}_{J}}\;{A_{1}}^{J}\;,\;\;\;\;\;J=0,8,$ (44)
where
$G_{S^{d}_{J}}=G_{S^{d}_{J}}(E_{A})$ $=$ $8\displaystyle 6\;d_{11J}\;\left[E_{\pi}(3G_{d2}^{R}-G_{d1}^{R})-E_{A}(3G_{d2}^{R}+G_{d1}^{R})\right],\;\;\;\;\;J=0,$ 8 . (45)
For the corresponding energy dependent decay rates of the axial vector meson decay into a positive or a negative pion and scalar, the following ratio can be defined:
$\mbox{Ratio}(ASP-A_{J})\;:\;\frac{\Gamma_{\pi^{+}}^{J}}{\Gamma_{\pi^{-}}^{J}}$ $\sim$ $\frac{\left|+G_{S^{d}_{J}}(P^{2})+G_{S^{f}_{J}}(P^{2})D_{A_{3}}(P^{2})\right|^{2}}{\left|+G_{S^{d}_{J}}(P^{2})-G_{S^{f}_{J}}(P^{2})D_{A_{3}}(P^{2})\right|^{2}},\;\;\;\;\;\;\;\;\;J=0,8.$ (46)
3.3 VPP: $\rho^{\pm}$ decay
Consider the interaction terms given in (16). By picking up only the f-channel terms describing the decay $\rho^{\pm}\to\pi^{\pm}\pi^{0}$, it yields:
${\cal L}_{\rho^{\pm}\pi^{0}\pi^{\pm}}^{f}$ $=$ $6\;i\;f_{123}\left[G_{d2}\left(-(\partial^{\mu}\rho^{+})\pi^{+}\eta+(\partial^{\mu}\rho^{-})\pi^{-}\eta\right)-\left(-\rho^{+}(\partial^{\mu}\pi^{+})\eta+\rho^{-}(\partial^{\mu}\pi^{-})\eta\right)\right.$ (47)
$+$ $\left.G_{d1}\left(-\rho^{+}\pi^{+}(\partial^{\mu}\eta)+\rho^{-}\pi^{-}(\partial^{\mu}\eta)\right)\right],$
The decay $\rho^{\pm}\to\pi^{\pm}\eta$ is described by the following terms:
${\cal L}_{\rho\pi\pi}^{f}$ $=$ $G_{V}^{f}\left[-\rho^{+}\pi^{+}\eta+\rho^{-}\pi^{-}\eta\right],$ (48)
where
$G_{V^{f}}=6\left[(E_{\rho}+E_{\pi^{\pm}})G_{d2}-E_{\pi^{0}}G_{d1}\right].$ (49)
As discussed above, in this work, the momentum dependence of the coupling constants are not taken into account. So, accordingly, meson energies in these equations are taken as their rest masses.
Interactions (16) also describe the decay $\rho^{\pm}\to\pi^{\pm}\eta$ in the channel with $T^{ijk}\to 2d^{ijk}$ that, with a mixing $\eta-\pi^{0}$, leads to the same final state as the decay (47),
${\cal L}_{\rho^{\pm}\eta\pi^{\pm}}^{d}$ $=$ $6\;i\;d_{110}\left[G_{d2}\left((\partial^{\mu}\rho^{+})\pi^{+}\eta+(\partial^{\mu}\rho^{-})\pi^{-}\eta\right)-\left(\rho^{+}(\partial^{\mu}\pi^{+})\eta+\rho^{-}(\partial^{\mu}\pi^{-})\eta\right)\right.$ (50)
$+$ $\left.G_{d1}\left(\rho^{+}\pi^{+}(\partial^{\mu}\eta+)+\rho^{-}\pi^{-}(\partial^{\mu}\eta)\right)\right].$
The mixing $\eta^{\prime}-\pi^{0}$ (channel $J=8$) is not considered because it is not energetically favorable, i.e. $M_{\eta^{\prime}}>M_{\rho}$.
By including the meson mixing in the d-channel, the amplitudes can be written as:
${\cal V}_{\rho^{\pm}\pi^{\pm}\pi}^{d}$ $=$ $G_{V^{d}}\;D_{\eta}(P^{2}_{\eta}),$ (51)
where
$G_{V^{d}}=6\sqrt{\frac{2}{3}}\;\bar{G}_{03}\;\left[G_{d2}(E_{\rho}+E_{\pi})-E_{\eta}G_{d1}\right].$ (52)
The $\eta-\pi^{0}$ mixing in the final state of the $\rho^{\pm}$ decay can be searched experimentally in spite of being a small effect.
The following ratio of decay rates for the charged rho can be extracted:
$\mbox{Ratio}(VPP)\;:\;\frac{\Gamma_{\pi^{+}}^{V}}{\Gamma_{\pi^{-}}^{V}}$ $\sim$ $\frac{\left|+G_{V^{d}}D_{\eta}(P^{2})-G_{V^{f}}\right|^{2}}{\left|+G_{V^{d}}D_{\eta}(P^{2})+G_{V^{f}}\right|^{2}}\;\;.$ (53)
3.3.1 VPA followed by VPP
The ratio of decay rates for the complete VPA-A followed by VPP decay - exhibited in Fig. (3) - can be written as:
$\mbox{Ratio}\;(VPA-VPPJ)\;:\;\frac{\Gamma_{\pi^{+}}^{J}}{\Gamma_{\pi^{-}}^{J}}$ $\sim$ $\frac{\left|\left[+G_{A_{J}}^{d}+G_{A_{J}}^{f}D_{A_{3}}(P^{2})\right]D_{\rho}(P_{\rho}^{2})\left[+G_{V^{d}}D_{\eta}(P^{2}_{\eta})-G_{V^{f}}\right]\right|^{2}}{\left|\left[+G_{A_{J}}^{d}-G_{A_{J}}^{f}D_{A_{3}}(P^{2})\right]D_{\rho}(P_{\rho}^{2})\left[+G_{V^{d}}D_{\eta}(P^{2}_{\eta})+G_{V^{f}}\right]\right|^{2}},$ (54)
where $J=0,8$, $D_{\rho}(P_{\rho}^{2})$ and $D_{\eta}(P_{\eta}^{2})$ are respectively the rho and eta propagators with corresponding momenta according to Figs. (3) and (2).
4 Numerical estimations
An effective transverse gluon propagator, extracted from Schwinger Dyson equations at the rainbow ladder approximation, was taken from Refs. [45] and used to compute the meson coupling constants:
$R_{T}(k)$ $=$ $\frac{8\pi^{2}}{\omega^{4}}De^{-k^{2}/\omega^{2}}+\frac{8\pi^{2}\gamma_{m}E(k^{2})}{\ln\left[\tau+(1+k^{2}/\Lambda^{2}_{QCD})^{2}\right]},$ (55)
with the following parameters $\gamma_{m}=12/(33-2N_{f})$, $N_{f}=4$, $\Lambda_{QCD}=0.234$GeV, $\tau=e^{2}-1$, $E(k^{2})=[1-exp(-k^{2}/[4m_{t}^{2}])/k^{2}$, $m_{t}=0.5GeV$, $\omega=0.5$GeV and $D=(0.55)^{3}/\omega$ (GeV2). It takes into account a quark-gluon running coupling constant and it was extracted from calculations with Schwinger Dyson equations at the rainbow ladder level.
The parameters used to perform numerical calculations are those from calculations for the (pseudoscalar) meson spectrum defined in [35]. The set of parameters $X-20$ $D_{I,2}$ of that work, calculated with the same gluon propagator of this work, provides the following values for the quark effective masses:
$M_{u}=0.392\;GeV,\;\;\;\;\;M_{d}=0.396\;GeV,\;\;\;\;\;M_{s}=0.600\;GeV,$ (56)
for $\Lambda=0.675$ Gev. These parameters are different from the parameters used in [7], and consequently resulting three-meson coupling constants are different (stronger). Concerning the mixing parameters $G_{i\neq j}$ (for $i,j=0,3,8$), whose overall normalizations are not unambiguous [35, 36], it is interesting to note they are directly proportional to (effective) quark mass differences. As a consequence, the hierarchies of mixings in all the channels pseudoscalar, scalar, vector and axial mesons, are usually the same, $G_{08}>G_{38}>G_{03}$, and this goes along with well known results [46]. To keep the normalization procedure consistent with the effective masses shown above, the following meson mixing parameters - for the same set of parameters $X-20$ $D_{I,2}$ of the quoted work - were adopted:
$G_{33}^{R}=10.00\;GeV^{-2},\;\;\;\;\;G_{88}^{R}=7.61\;GeV^{-2},\;\;\;\;\;G_{00}^{R}=8.60\;GeV^{-2},$
$G_{03}^{R}=0.04\;GeV^{-2},\;\;\;\;\;\;G_{08}^{R}=1.80\;GeV^{-2},\;\;\;\;\;\;G_{38}^{R}=0.05\;GeV^{-2}.$ (57)
It follows from Eq. (33) that:
$\bar{G}_{03}=0.0011,\;\;\;\;\;\;\bar{G}_{08}=0.0555,\;\;\;\;\;\;\bar{G}_{38}=0.0014.$ (58)
There are corrections to these values from the sixth order quark interactions [7] which are usually smaller and will not be taken into account for the present estimations.
The resulting values of the three-meson coupling constants used in this work are the following:
$G_{sb,1}^{R}=-0.216\;GeV,\;\;\;\;\;\;G_{d1}^{R}=-0.551,\;\;\;\;\;G_{d2}^{R}=-0.150,\;\;\;\;\;G_{sva,sb}^{R}=0.018\;GeV^{-1}.$ (59)
[…]
In Fig, (4) the Ratios VPA-A, Eq. (37), for $J=0$ and $J=8$ are shown as function of the energy $P^{2}$. There are interference effects that drive the ratio of decay rates $\Gamma_{\pi^{+}}/\Gamma_{\pi^{-}}$ to values smaller than one, i.e. destructive interference for positive pion production and constructive interference for negative pion production. This is rather noted in the range of energies around the resonance mass of the intermediary states $a_{1}(1260)$ induced by meson mixing, being $M_{A_{1}}^{2}\simeq 1.6$ GeV2. The symbols for the ratio are different below and above the physical threshold $P^{2}=M_{A_{0}}^{2}$ (where $A=f_{1}(1285)$ and $f_{1S}(1420)$). Maybe there is a way to verify experimentally if there is a threshold (and which would be) for the corresponding different production rates of positive and negative pions. The non-zero width of the intermediary state also plays important roles in the resulting behavior. The ratio tends to one for energies larger than $\sim 4-5$ GeV2. The production of positive and negative rhos, however, has the opposite asymmetry and it contains an excess of positive rhos in these VPA channel. Since the preferential decay of such positive and negative rhos is $\rho^{\pm}\to\pi^{0}\pi^{\pm}$, one may expect that these decays from $\rho^{\pm}$ tend to restore the ratio $\Gamma_{\pi^{+}}/\Gamma_{\pi^{-}}=1$ in this mechanism. However, it may happen that other processes, such as inelastic scattering or absorption of charged rhos, reduce the number of charged rhos available in the (strongly interacting) final state. Consequently, it is possible to expect that the ratio of charged pions from the VPA channel of axial meson decays is different from the one shown the figure (4). If positive rhos are suppressed in this final state, more than negative rhos, this ratio would reach still smaller values. However, there is a scenario in which negative rhos are absorbed more often than positive rhos, in which case the ratio of positive to negative produced pions would be larger than one, in this channel. Another possibility, envisaged below, is that the charged rho decay also presents interference effects.
Figure 4: Ratio of decay rates of Eq. (37) (VPA) for $f_{1}(1285)$ and $f_{1S}(1420)$ (J=0, 8).
The ratio (46) for the decay with interaction ASP is exhibited for $J=0,8$ in Fig. (5). The ratio has the same behavior as the VPA channel shown above with the same range of energies, exhibiting destructive (constructive) interference for production of positive (negative) pions. The interference strength however is smaller due to the relative values of the involved coupling constants. The behavior with energy is the same and the ratio go to one around $P^{2}=3.5$ GeV2 or $P^{2}=5.0$ GeV2 depending on the channel. A similar effect to one discussed for the channel VPA may arise due to the fact that $a_{0}^{\pm}$ may decay into $\pi^{\pm}\eta$ [25] and this could restore the same number of positive and negative pions in the final state. However, these charged scalar mesons have a width smaller than the width of the rho meson and they may be expected to undergo different effects or scattering in the final state that would prevent them to decay into charged pions. The final difference between the number of positive and negative pions, from this mechanism, may be therefore smaller than the one indicated in Fig. (5). Furthermore, the energy dependencies of the couplings - Eqs. (7) and (39) - lead to decreasing functions of the energy and these effects may lead to smaller asymmetry in the ratio. Similarly to the VPA channel discussed above, several scenarios are possible.
Figure 5: Ratio of decay ratios of Eq. (46) (ASP-A) processes.
In Fig. (6) the ratio (53) is presented as a function of the energy $P^{2}\equiv P^{2}_{\eta}$, by using different symbols above and below the rho rest energy. The maximum interference takes place around $P^{2}\sim M_{\eta}^{2}$ being the width of the $\eta$ very small. The small deviation from 1 is due to the relative values of the coupling constants. The complete energy dependence of the interactions should make these values still smaller. By considering the initial state $\rho^{\pm}$ in its rest frame, the following CP breaking parameter have been calculated
$A^{\rho}$ $=$ $\frac{\Gamma_{\pi^{+}}^{V}-\Gamma_{\pi^{-}}^{V}}{\Gamma_{\pi^{+}}^{V}+\Gamma_{\pi^{-}}^{V}},$ (60)
where
$\Gamma_{\rho^{\pm}}^{V}$ $=$ $\frac{|M_{VPP,\pi^{\pm}}|^{2}|\vec{K}_{\pi}|}{8\pi M_{\rho}^{2}},$ (61)
where $|\vec{K}_{\pi}|$ s the pion momentum and $M_{VPP}$ is the amplitude for the corresponding channel written in Eqs. (52), (49) and (53). The charged pion momentum is not the same in the d-channel (where it shares energy with the $\eta$ from the VPP decay) and in the f-channel. The corresponding pion momentum was considered for each of the amplitudes of the d-channel and f-channel, being that, in the interference term of $|M_{VPP,\pi^{\pm}}|^{2}$, an average value was taken. For the charged rho decays the resulting value is:
$A^{\rho}\simeq 2.5\times 10^{-4},$ (62)
This parameter is quite small and it might be reduced by the momentum and energy dependence of the interactions or if the $\eta-\pi^{0}$ mixing is suppressed for some reason. It is interesting to emphasize however, that the interference takes place below the threshold of $M_{\rho}$ as seen in Fig. (6).
Figure 6: Ratio of decay ratios of Eq. (53) (VPP) for $P^{2}\equiv P_{\eta}^{2}$, different symbols above and below the threshold of the rho rest energy $P^{2}=M_{\rho}^{2}$.
The ratio of the decay rate (54) for the VPA interaction with subsequent VPP decay of the rho meson, by including the possible $\eta-\pi^{0}$ that leads to interference, as shown in Fig. (6), is exhibited in Fig. (7). In this case, the d-channel and the f-channel of the VPP decay are possible in both d- and f-channels of the VPA decay. Different symbols are used for energies below and above the threshold of each case, $f_{1}(1285)$ and $f_{1S}(1420)$. The resulting asymmetry is smaller than in the VPA case without VPP decay presented in Fig. (4).
Figure 7: Ratio of decay ratios of Eq. (54) for VPA-A process with the second decay VPP for both cases, decays of $f_{1}(1285)$ (J=0) and of $f_{1S}(1420)$ (J=8).
As discussed above, and in [35], the normalization of the meson mixing parameters are not unambiguous. In Fig. (8) the ratios VPA and ASP, Eqs. (37) and (46), are presented by considering, arbitrarily, five times larger mixing interactions than those used above: $5\times\bar{G}_{i\neq j}$ for $i,j=0,3,8$. Whereas the behavior of the ratio remains the same, by comparing with results of the figures above, it is seen that the interference effect is amplified proportionally to the strength of the mixing interactions.
Figure 8: Ratios of decay ratios of decay to positive to negative pions for VPA and ASP - 0 and 8 - processes, by considering mixing interactions five times larger than the ones used in the previous figures.
In Fig. (9) the ratios of Eqs. (37) and (46), for VAP and ASP channels (with $J=0$ and $8$), are exhibited for the mixing interactions with opposite sign. The constructive and destructive interference in the production of negative and positive pions are exchanged.
Figure 9: Ratios of decay ratios of decay to positive and negative pions, for VAP and ASP with J= 0 and 8 processes, Eqs. (37) and (46), by considering mixing interactions with signal minus, $-\bar{G}_{i\neq j}$ for i,j=0.3,8.
5 Summary
Decays of two light neutral axial mesons, $f_{1}(1285)$ and $f_{1S}(1420)$, into final states with charged pions were investigated in this work: $\;A\to\rho^{\mp}+\pi^{\pm}$ (VPA), and $\;A\to a_{0}^{\mp}+\pi^{\pm}$ (ASP). These processes are energetically favorable contrarily to other possible channels of the same three-meson interactions. The $f_{1}(1285)$ decays considered in this work have already been observed experimentally [25] and the decays of $f_{1S}(1420)$ have not. The reasons for this lack of experimental observation are not clear, being possibly related to very small branching ratio, difficulties with final state interactions and decays or even that its structure does not correspond exactly to the state $A_{8}$ of the quark-antiquark nonet. However, the state $A_{8}$ of the possible axial meson nonet can undergo strong mixings with the states $A_{0}$ and $A_{3}$, as discussed in Section (3.1.1), and these mixings can prevent a clear identification of its structure at the example of the $\eta-\eta^{\prime}$ mixing of the pseudoscalar meson multiplet. Therefore, this investigated was done as a possible scenario. The resulting asymmetry in positive and negative produced pion, due to interference, is small, and it is directly proportional to the strength of the neutral meson mixing parameters for $f_{1}(1285)-a_{1}(1260)$ and $f_{1S}(1420)-a_{1}(1260)$. The maximum of the interference occurs below the threshold energy for the on shell axial meson, either $f_{1}(1285)$ or $f_{1S}(1420)$, around the rest energy of the intermediary state $a_{1}(1260)$, $M_{a_{1}}$. Experimental results concerning axial meson mixings may be very difficult to be obtained and they are missing. This mechanism of asymmetry in the production of charged pions due to interference is suppressed at high energies. There are at least two effects that might reduce further these values, or could even change their signs in some scenarios. The complete resulting momentum dependencies of the used coupling constants are decreasing functions for higher energies, and this decrease is faster than the possible energy dependencies of the interactions in (6) and (9). Secondly, the subsequent decay of charged rhos or scalars $a_{0}^{0}$ into charged pions may reduce these values if these charged particles ($\rho^{\pm}$ and $a_{0}^{\pm}$) do not undergo other inelastic process that prevents than to decay into a charged pion. Subsequent final state interaction with absorption or inelastic scattering of $\rho^{\pm}$ and/or $a_{0}^{\pm}(980)$ would play an important role in the resulting ratios. Therefore, further investigation is needed.
The decay $\rho^{\pm}\to\pi^{\pm}\pi^{0}$ was also investigated with corresponding three-meson vertices. The interference is relatively smaller than in the decay of the axial mesons and, for the interactions used, it is only possible if the mixing $\eta-\pi^{0}$ is considered. This mixing may be tested in decays of $\rho^{\pm}$. However, for the present mechanism, the mixing happens below the threshold of the rho rest energy. A CP-breaking parameter for this decay was estimated to be $A^{\rho}\simeq 2.5\times 10^{-4}$. Information from experimental observation seem missing. This decay was also considered for the amplitude of the VPA process, in which the decay of the resulting rho also presents interference. There is a slight decrease in the overall interference, and therefore in the ratio $\Gamma_{\pi^{+}}/\Gamma_{\pi^{-}}$.
Therefore, the interference mechanism of asymmetry in charged pion production can be related to controversial problems: neutral meson mixings; scalar and axial meson structures; and the behavior of the final states, i.e, the possibility of different absorption rates or mechanisms for charged $\rho^{\pm}$ and $a_{0}^{\pm}(980)$. It was assumed that the scalar and axial mesons fit, at least in part, into a flavor U(3) nonet although this is not an uncontroversial subject [23, 25, 24, 47]. Several other types of interactions among pseudoscalar, scalar, axial and vector mesons/states, involving different meson states with strangeness, within this kind of processes, were left outside the scope of the work. In addition to that, a more complete account of energy dependent effects with more complete energy and momentum dependencies are currently investigated and will be presented in another work.
Acknowledgements
F.L.B. is a member of INCT-FNA, Proc. 464898/2014-5. F.L.B. acknowledges partial support from CNPq-312750/2021 and CNPq-407162/2023-2.